Core refractive index -1.55
Cladding refractive index is 1.53wht is numerical aperture for optical fibre
1
Expert's answer
2020-11-25T10:46:37-0500

As per the given question,

the refractive index of the core "(n_1)=1.55"

Refractive index of cladding "(n_2)=1.53"

Numerical aperture of the optical fiber"(NA)=\sqrt{n_1^2-n_2^2}"

Now, substituting the values in the above,

"NA=\sqrt{1.55^2-1.53^2}"

"\Rightarrow NA=\sqrt{0.0616}"

"\Rightarrow NA=0.248\sim0.25"